FIND the LCM of the set of polynomials: t^2-t-12 and t^2+2t-24.
Thanks

Answer by venugopalramana(3286)

FIND the LCM of the set of polynomials: t^2-t-12
=T^2-4T+3T-12=T(T-4)+3(T-4)=(T+3)(T-4).......................I
and t^2+2t-24.
=T^2+6T-4T-24=T(T+6)-4(T+6)=(T-4)(T+6)..............................II
(T-4) IS A COMMON FACTOR.
HENCE LCM IS
(T-4)(T+3)(T+6)